OpenStudy (anonymous):

Two congruent "door-stop" trapezoids are placed edge-to-edge with measurements and markings as shown. Find the length of segment AC. I cant figure out how to find the answer.

OpenStudy (anonymous):

this is the figure

OpenStudy (anonymous):

use pythagoras BE^2=25^2-20^2=625-400=225 sq rt.=15 so BE=15 BE=BC=15=AB given so AC=30